# Perfect t-embeddings of doubly periodic Aztec diamonds

### v1

@@ 1. Introduction
A dimer configuration on a graph is a subset of edges covering every vertex exactly once. Given edge weights, the dimer model is a probability measure on the set of dimer configurations, where each configuration has probability proportional to the product of its edge weights. The dimer model on weighted planar bipartite graphs is known to exhibit a rich set of behaviors. One of its most striking features is the existence of three distinct phases – frozen, liquid, and gaseous (also known as frozen, rough, and smooth) – within a single system. In the frozen phase, the dimer configuration becomes rigid and highly ordered, with no local fluctuations; certain patterns dominate, and entropy is minimal. In contrast, the liquid phase exhibits long-range correlations, with the variance of the height fluctuations growing logarithmically. Meanwhile, in the gaseous phase, the height fluctuations have bounded variance. These phases often coexist within a single large configuration, separated by sharp interfaces called arctic curves, leading to a striking interplay of order and randomness that makes the dimer model a central object in statistical mechanics and combinatorics. In recent years, systems exhibiting all three phases simultaneously have attracted considerable attention, as evidenced by the following (non-exhaustive) list of recent works, [4, 10, 11, 12, 30, 29, 8, 9, 39, 6, 5, 33, 56, 28, 19, 13, 18, 20, 48, 14, 31, 32, 49, 55, 57, 52, 46].
@@
Recall that a t-embedding is a proper embedding of the dual graph in which the dual edge lengths define the same probability measure as the original edge weights, and at each interior vertex, the angles around adjacent white and black faces both sum to $\pi$. Together with a t-embedding, we consider the corresponding origami map of the dual graph. Informally speaking, to get the origami map out of the t-embedding, one should fold the plane along each edge of the t-embedding. The angle condition guarantees that this folding procedure is consistent.
@@
One of the natural questions is how the limiting t-surfaces are influenced by the emergence of frozen and gaseous regions. In this paper, we investigate this question in the context of a rich and structured example: the doubly periodic dimer model on the Aztec diamond. In previously studied settings, such as the uniform dimer measure on the Aztec diamond [27, 16], the tower graph [16], and the hexagon [15] – it is known that: each frozen region collapses to a point on the boundary under both t-embedding and origami map; the limit of the origami map lies on a straight line; the image of the liquid region under the t-surface converges to a maximal surface in Minkowski space $\mathbb{R}^{2,1}$.
@@
In this paper, we study the limiting t-surface of the Aztec diamond with $(k\times\ell)$-periodic weights (i.e., edge weights that are periodic in both coordinate directions). It is known that when both the parameters $k$ and $\ell$ are at least two, all three types of macroscopic regions might appear in the limit of large domains – frozen, liquid, and gas regions. We focus on two specific subfamilies of weights, analyzing each case separately. The first family corresponds to the $(1\times\ell)$-periodic case (with some mild additional assumptions), which features $2(\ell+1)$ frozen regions and no gas regions. The second family is within the $(2\times\ell)$-periodic setting, exhibiting $4$ frozen regions and $\ell-1$ gas regions. This second family coincides with the model studied in [31, 10]. We focus on this setting because it is rich enough to exhibit arbitrarily many gas regions, while retaining sufficient symmetry to simplify our analysis. In particular, these symmetries allow us to rely on both the rather elementary techniques used in [10] as well as the more technical machinery developed in [11].
@@ 1.1.2. Aztec diamonds with ($2\times\ell$)-periodic weights
Similarly to (1.3), we first define the limiting functions for the t-embedding and its origami map on the corresponding spectral curve. For $q=(z,w)\in\mathcal{R}$ define
@@
where $Q$ is a matrix mentioned above and defined by (2.16), and $\gamma_{q}$ is a symmetric under conjugation simple curve in $\mathcal{R}$ going from $q$ to $\bar{q}$ precisely described in Definition 5.12. As in Section 1.1.1 above, for simplicity we use continuum coordinates (5.29) to index faces in the Aztec diamond.
@@
Let $\mathcal{T}_{N}$ and $\mathcal{O}_{N}$ denote the perfect t-embedding and its associated origami map of the Aztec diamond of size $2\ell N$ with $(2\times\ell)$-periodic weights described above. Let $(\xi,\eta)$ be a point in the liquid region $\mathcal{F}_{R}$ of the Aztec diamond. Then
as $N\to\infty$, where $\Omega:\mathcal{F}_{R}\to\mathcal{R}_{0}$ is a diffeomorphism (properly defined in Proposition 5.10) known as the critical point map, and the convergence is uniform on compact subsets of $\mathcal{F}_{R}$.
Moreover, let $(\xi,\eta)$ be in a gas or frozen region, and let $(\xi^{\prime},\eta^{\prime})\in\mathcal{F}_{R}$ be a sequence in the liquid region converging to a point $(\xi_{0},\eta_{0})$ on the boundary of the liquid region and the gas or frozen component containing $(\xi,\eta)$, then
$\lim_{N\to\infty}(\mathcal{T}_{N}(\xi,\eta),\mathcal{O}_{N}(\xi,\eta))=\lim_{(\xi^{\prime},\eta^{\prime})\to(\xi_{0},\eta_{0})}\left(\mathcal{Z}(\Omega(\xi^{\prime},\eta^{\prime})),\vartheta(\Omega(\xi^{\prime},\eta^{\prime}))\right).$ (1.8)
@@
Note that part of the statement is that the right-hand side of (1.8) depends only on the gas or frozen connected component containing $(\xi,\eta)$, and not on the specific choice of $(\xi_{0},\eta_{0})$.
See Theorem 5.15 and Corollary 5.16 for the precise statements. The proof follows a similar strategy as the proof of Theorem 1.1. We employ Theorem 1.14 and the formulas for the inverse Kasteleyn matrix from [10], and after some computations we obtain an exact formula for $\mathcal{T}_{N}$ and $\mathcal{O}_{N}$ in Theorem 5.5. The limits are then obtained from a steepest descent analysis using the technique developed in [11]. In fact, the action function is the same as in [10, 11], which allow us to re-use their arguments almost word by word.
@@
Recall that in our setup there are $g=\ell-1$ gas components. Let us enumerate the gas components by $i=1,\dots,\ell-1$. We denote the right hand side of (1.8) by $P_{i}\in\mathbb{R}^{2,2}$, if $(\xi_{0},\eta_{0})$ is in the $i$th gas region. We are now ready to describe the geometry of the scaling limit of the t-surface. See Theorem 5.26 in the text.
Theorem 1.9.
1) The t-surface $(\mathcal{T}_{N},\mathcal{O}_{N})$ of the Aztec diamond of size $2\ell N$ with $(2\times\ell)$-periodic weights (as shown in Figure 4) converges to a space-like surface in $\mathbb{R}^{2,2}$ with zero mean curvature and with the boundary given by a quadrilateral in $\mathbb{C}\times\mathbb{R}$ with vertices at
and $\ell-1$ cusps with apices $P_{1},\dots,P_{\ell-1}$;
2) In the scaling limit, each of the $4$ frozen regions collapses to one of the boundary points, and each of the $(\ell-1)$ gas regions collapses to one of the inner points $P_{i}$ under the t-surface.
@@ 1.1.3. Coulomb gauges for general graphs with outer face of degree four
It is known [42, Theorem 2] that a perfect t-embedding of a finite planar bipartite graph $\mathcal{G}$ with the outer boundary of degree four and real-valued Kasteleyn weights $K_{\mathcal{G}}$ always exists. This follows from the fact that t-embeddings of $\mathcal{G}^{*}$ are preserved under elementary transformations of $\mathcal{G}$ and $\mathcal{G}$ can be reduced to the 4-cycle graph by applying a sequence of elementary transformations, without modifying the 4 boundary vertices at intermediate stages. This gives a construction of a perfect t-embedding using a sequence of elementary transformations simplifying the graph to a 4-cycle.
@@ 1.1.4. Outlook
The paper is organized as follows. Section 2 contains definitions and statements of relevant facts about both perfect t-embeddings and the spectral curve of the dimer model in the settings we consider. In Section 3 we present and prove formulas for Coulomb gauge functions in terms of the inverse Kasteleyn matrix for general planar weighted bipartite graphs with outer face of degree four. In Section 4 we specialize this formula to the $(1\times\ell)$-periodic case and take asymptotics, proving Theorem 1.1 and Corollary 1.2. Section 5 is devoted to the $(2\times\ell)$-periodic case. In this section, we prove Theorem 1.6, Theorem 1.9 and Proposition 1.12; in particular we analyze the cusps which form in the interior of the t-surface. In Section 6 we specialize the weights to what is known as the two-periodic Aztec diamond, and analyze how the cusp changes when the weights in the fundamental domain are simply shifted.
@@ 2.4. The spectral data
Being standard means that $e_{v}\in\mathbb{R}^{g}/\mathbb{Z}^{g}$. These divisors are naturally described through the discrete Abel map $\mathbf{d}$, introduce in [36, Section 3], and a parameter $t\in\mathbb{R}^{g}/\mathbb{Z}^{g}$. Namely, for each white vertex $w$ and each black vertex $b$,
@@
The discrete Abel map is defined on the vertices of the quad graph, that is, on the union of the vertices in the fundamental domain and its dual, and is defined through the image of the angles under the Abel map. We refer to [22, Section 3.2] for a definition and many properties and to [11, Section 5.4] for a discussion closer to our setting. See also [22, Remark 50] for a discussion on the parameter $t$.
@@
Both the theta function and the prime form can be used as building blocks of meromorphic functions and forms on the Riemann surface, in a similar spirit as linear terms $az+b$ are building blocks for meromorphic functions on the Riemann sphere. There is a remarkable identity, known as Fay’s identity, that involves both the theta functions and prime forms. The version that will be relevant for us is the following, see [35, Proposition 2.10]. For $q,p_{1},p_{2}\in\mathcal{R}$ and $e\in\mathbb{C}^{g}$,
@@ 6. The two-periodic Aztec diamond
The second part of the statement follows by taking the integrals over the curve $\gamma_{q^{\prime}}$ where $q^{\prime}\in A_{1}$}. By the reciprocity law (see [35, Equation (7)]) or, from Riemann bilinear relation (see [34, Section III.3]). We have

### v2

@@ 1. Introduction
A dimer configuration on a graph is a subset of edges covering every vertex exactly once. Given edge weights, the dimer model is a probability measure on the set of dimer configurations, where each configuration has probability proportional to the product of its edge weights. Asymptotic behavior of the model is often characterized via the height function associated to each dimer configuration, first defined by Thurston [59]. The dimer model on weighted planar bipartite graphs is known to exhibit a rich set of behaviors. One of its most striking features is the existence of three distinct phases – frozen, liquid, and gaseous (also known as frozen, rough, and smooth) – within a single system. In the frozen phase, the dimer configuration becomes rigid and highly ordered, with no local fluctuations; certain patterns dominate, and entropy is minimal. In contrast, the liquid phase exhibits long-range correlations, with the variance of the height fluctuations growing logarithmically. Meanwhile, in the gaseous phase, the height fluctuations have bounded variance. These phases often coexist within a single large configuration, separated by sharp interfaces called arctic curves, leading to a striking interplay of order and randomness that makes the dimer model a central object in statistical mechanics and combinatorics. In recent years, systems exhibiting all three phases simultaneously have attracted considerable attention, as evidenced by the following (non-exhaustive) list of recent works, [4, 10, 11, 12, 31, 30, 8, 9, 40, 6, 5, 34, 57, 29, 19, 13, 18, 22, 49, 14, 32, 33, 50, 56, 58, 53, 47].
@@
Recall that a t-embedding is a proper embedding of the augmented dual graph in which the dual edge lengths define the same probability measure as the original edge weights, and at each interior vertex, the angles around adjacent white and black faces both sum to $\pi$. Together with a t-embedding, we consider the corresponding origami map of the dual graph. Informally speaking, to get the origami map out of the t-embedding, one should fold the plane along each edge of the t-embedding. The angle condition guarantees that this folding procedure is consistent. A t-embedding is said to be a perfect t-embedding if its boundary polygon is tangential to a circle, and if edges connecting outer vertices with interior ones lie on angle bisectors.
@@
In previously studied settings, such as the uniform dimer measure on the Aztec diamond [28, 16], the tower graph [16], and the hexagon [15], it is known that: each frozen region collapses to a vertex of the boundary polygon under both the perfect t-embedding and its origami map; the limit of the origami map lies on a straight line; the image of the liquid region under the t-surface converges to a maximal surface in Minkowski space $\mathbb{R}^{2,1}$. An interesting open question is how the emergence of quasi-frozen and gaseous regions affects the limiting t-surfaces. This paper provides the first investigation of this question, focusing on a rich and highly structured example: the doubly periodic dimer model on the Aztec diamond.
It is conjectured that for dimer models exhibiting gaseous phases, the t-surface converges to a maximal surface with cusps, while it is an open question whether the ambient space may enlarge from $\mathbb{R}^{2,1}$ to $\mathbb{R}^{2,2}$, that is, whether the origami map is one- or two-dimensional in the limit. Our main theorem confirms this conjecture and settles the open question. In addition, it shows that, as in the previously studied examples, the t-embedding identifies the conformal structure governing the scaling limit of height fluctuations: in the case without gaseous regions, the Gaussian Free Field, and in the case with gaseous regions, the mixture of the Gaussian Free Field with a discrete Gaussian. Informally, our main result is the following; see Section 1.1, and in particular Theorem 1.6 and Proposition 1.12, for precise statements.
Theorem.
The scaling limit of the t-surface is a space-like maximal surface with light-like cusps which are in one-to-one correspondence with the gaseous facets. In general, the surface is contained in $\mathbb{R}^{2,2}$ and not in $\mathbb{R}^{2,1}$. Moreover, the conformal structure on the liquid region defined from the maximal surface coincides with the Kenyon–Okounkov conformal structure.
@@
In slightly more detail, we study the limiting t-surface of the Aztec diamond with $(k\times\ell)$-periodic weights (i.e., edge weights that are periodic in both coordinate directions). It is known that when both the parameters $k$ and $\ell$ are at least two, all three types of macroscopic regions might appear in the limit of large domains – frozen, liquid, and gas regions. We focus on two specific subfamilies of weights, analyzing each case separately. The first family corresponds to the $(1\times\ell)$-periodic case (with some mild additional assumptions), which features $2(\ell+1)$ frozen regions and no gas regions. The second family is within the $(2\times\ell)$-periodic setting, exhibiting $4$ frozen regions and $\ell-1$ gas regions. This second family coincides with the model studied in [32, 10]. We focus on this setting because it is rich enough to exhibit arbitrarily many gas regions, while retaining sufficient symmetry to simplify our analysis. In particular, these symmetries allow us to rely on both the rather elementary techniques used in [10] as well as the more technical machinery developed in [11].
@@ 1.1.2. Aztec diamonds with ($2\times\ell$)-periodic weights
Similarly to (1.3), we first define the limiting functions for the t-embedding and its origami map on the corresponding spectral curve. For $q=(z,w)\in\mathcal{R}_{0}$ define
@@
where $Q$ is a matrix mentioned above and defined by (2.16), and $\gamma_{q}$ is a symmetric under conjugation simple curve in $\mathcal{R}$ going from $q$ to $\bar{q}$ precisely described in Definition 5.12.
With these definitions, we move on to state the main theorem of this work, which includes three closely related statements. The first part is an explicit computation of the scaling limit of $\mathcal{T}_{N}$ and $\mathcal{O}_{N}$, the perfect t-embedding and origami map in the $(2\times\ell)$-periodic setting, in terms of the maps $\mathcal{Z}$ and $\vartheta$. The second statement says that gas regions collapse to interior points in the quadrilateral containing the embedding, and frozen regions collapse to boundary points. The third statement says that the image of the map defined by the pair $(\mathcal{Z},\vartheta)$ is a space-like maximal surface in $\mathbb{R}^{2,2}$, with cusps which are in one-to-one correspondence with the gaseous facets.
Let $(u,v)=(2(\ell x+i)+\varepsilon,2(2y+j)+\varepsilon)$, with $x,y\in\mathbb{Z}$, $i=0,\dots,\ell-1$, $j=0,1$, and $\varepsilon=0,1$, denote the coordinates of the faces in the Aztec diamond (see Section 2.2 and Figure 7). Let $(\xi,\eta)\in[-1,1]^{2}$ be the global coordinates in the scaled Aztec diamond, related to $(u,v)$ via (5.29).
@@
Let $\mathcal{T}_{N}$ and $\mathcal{O}_{N}$ denote the perfect t-embedding and associated origami map of the Aztec diamond of size $2\ell N$ with $(2\times\ell)$-periodic weights described above and let $(\xi,\eta)$ and $(u,v)$ be as above.
1) If $(\xi,\eta)$ is a point in the liquid region $\mathcal{F}_{R}$, then
$(\mathcal{T}_{N}(u,v)),\mathcal{O}_{N}(u,v)))\to\left(\mathcal{Z}(\Omega(\xi,\eta)),\vartheta(\Omega(\xi,\eta))\right)$
as $N\to\infty$, where $\Omega:\mathcal{F}_{R}\to\mathcal{R}_{0}$ is a diffeomorphism (precisely defined in Proposition 5.10) known as the critical point map, and the convergence is uniform on compact subsets of $\mathcal{F}_{R}$;
2) If $(\xi,\eta)$ is in a gas or frozen region, let $(\xi^{\prime},\eta^{\prime})\in\mathcal{F}_{R}$ be a sequence in the liquid region converging to a point $(\xi_{0},\eta_{0})$ on the boundary of the liquid region and the gas or frozen component containing $(\xi,\eta)$, then
$\lim_{N\to\infty}(\mathcal{T}_{N}(u,v),\mathcal{O}_{N}(u,v))=\lim_{(\xi^{\prime},\eta^{\prime})\to(\xi_{0},\eta_{0})}\left(\mathcal{Z}(\Omega(\xi^{\prime},\eta^{\prime})),\vartheta(\Omega(\xi^{\prime},\eta^{\prime}))\right);$ (1.8)
3) The map $\mathcal{R}_{0}\ni q\mapsto(\mathcal{Z}(q),\vartheta(q))$ defines a conformal parameterization of a space-like surface in $\mathbb{R}^{2,2}$ with zero mean curvature and with the boundary given by a quadrilateral in $\mathbb{C}\times\mathbb{R}$ with vertices at
and $\ell-1$ cusps with apices $P_{1},\dots,P_{\ell-1}$, where $P_{i}$ is the right hand side of (1.8) with $(\xi_{0},\eta_{0})$ in the boundary of the $i$th gas region.
@@
Part of the statement is that the right-hand side of (1.8) depends only on the gas or frozen connected component containing $(\xi,\eta)$, and not on the specific choice of $(\xi_{0},\eta_{0})$.
Remark 1.8.
One can obtain a conformal structure on the liquid region by pulling back the conformal structure on the surface defined by the limit of $(\mathcal{T}_{N},\mathcal{O}_{N})$ (described in the theorem above). Parts 1) and 3) of the theorem imply that this conformal structure matches the conformal structure obtained by pulling back the conformal structure of $\mathcal{R}_{0}$ via the map $\Omega$, which is known to coincide with the Kenyon-Okounkov conformal structure.
See Theorems 5.155.26 and Corollary 5.16 for the precise statements. The proof follows a similar strategy as the proofs of Theorem 1.1 and Corollary 1.2. We employ Theorem 1.14 and the formulas for the inverse Kasteleyn matrix from [10], and after detailed computations we obtain an exact formula for $\mathcal{T}_{N}$ and $\mathcal{O}_{N}$ in Theorem 5.5. The limits are then obtained from a steepest descent analysis using the technique developed in [11]. In fact, the action function is the same as in [10, 11], which allow us to re-use their arguments almost word by word. The proof of the third part of the theorem is provided in Section 5.6. It follows the strategy developed in [27, 16, 15], however, the arguments must be adapted to deal with the fact that the liquid region is multiply connected.
@@ 1.1.3. Coulomb gauges for general graphs with outer face of degree four
It is known [43, Theorem 2] that a perfect t-embedding of a finite planar bipartite graph $\mathcal{G}$ with the outer boundary of degree four and real-valued Kasteleyn weights $K_{\mathcal{G}}$ always exists. This follows from the fact that t-embeddings of $\mathcal{G}^{*}$ are preserved under elementary transformations of $\mathcal{G}$ and $\mathcal{G}$ can be reduced to the 4-cycle graph by applying a sequence of elementary transformations, without modifying the 4 boundary vertices at intermediate stages. This gives a construction of a perfect t-embedding using a sequence of elementary transformations simplifying the graph to a 4-cycle.
@@ 1.1.4. Outlook
The paper is organized as follows. Section 2 contains definitions and statements of relevant facts about both perfect t-embeddings and the spectral curve of the dimer model in the settings we consider. In Section 3 we present and prove formulas for Coulomb gauge functions in terms of the inverse Kasteleyn matrix for general planar weighted bipartite graphs with outer face of degree four. In Section 4 we specialize this formula to the $(1\times\ell)$-periodic case and take asymptotics, proving Theorem 1.1 and Corollary 1.2. Section 5 is devoted to the $(2\times\ell)$-periodic case. In this section, we prove Theorem 1.6 and Proposition 1.12; in particular we analyze the cusps which form in the interior of the t-surface. In Section 6 we specialize the weights to what is known as the two-periodic Aztec diamond, and analyze how the cusp changes when the weights in the fundamental domain are simply shifted.
@@ 2.4. The spectral data
Being standard means that $e_{v}\in\mathbb{R}^{g}/\mathbb{Z}^{g}$. These divisors are naturally described through the discrete Abel map $\mathbf{d}$, introduce in [37, Section 3], and a parameter $t\in\mathbb{R}^{g}/\mathbb{Z}^{g}$. Namely, for each white vertex $w$ and each black vertex $b$,
@@
The discrete Abel map is defined on the vertices of the quad graph, that is, on the union of the vertices in the fundamental domain and its dual, and is defined through the image of the angles under the Abel map. We refer to [21, Section 3.2] for a definition and many properties and to [11, Section 5.4] for a discussion closer to our setting. See also [21, Remark 50] for a discussion on the parameter $t$.
@@
Both the theta function and the prime form can be used as building blocks of meromorphic functions and forms on the Riemann surface, in a similar spirit as linear terms $az+b$ are building blocks for meromorphic functions on the Riemann sphere. There is a remarkable identity, known as Fay’s identity, that involves both the theta functions and prime forms. The version that will be relevant for us is the following, see [36, Proposition 2.10]. For $q,p_{1},p_{2}\in\mathcal{R}$ and $e\in\mathbb{C}^{g}$,
@@ 6. The two-periodic Aztec diamond
The second part of the statement follows by taking the integrals over the curve $\gamma_{q^{\prime}}$ where $q^{\prime}\in A_{1}$}. By the reciprocity law (see [36, Equation (7)]) or, from Riemann bilinear relation (see [35, Section III.3]). We have